OpenStudy (anonymous):

Explain how to get the product of 2 times the square root of 5 and 4 times the square root of 10

jhonyy9 (jhonyy9):

(2sqrt5)(4sqrt10)=8sqrt50=8sqrt(2*25)=8*5sqrt2=40sqrt2

jhonyy9 (jhonyy9):

ok ?

jhonyy9 (jhonyy9):

multiply 2 by 4 and those terms inside squarroot so 5 by 10